OpenStudy (anonymous):

Find S25 for 3 + 7 + 11 + 15 + …

OpenStudy (taylorterminate):

1] d = 7 - 3 = 4 a = 3 Sn = n/2 [2a + (n-1)d] S25 = 25/2 [2*3 + (25-1)4] = 25/2 [6 + 96] = 25 * (3 + 48) = 25 * 51 = 1275 2] d = 1/2 - 0 = 1/2 3] There are 4 terms between 40 and 10. d = (10 - 40) / (4+1) = -30/5 = -6 a = 40 AP is 40, 34, 28, 22, 16, 10 4] d = 0 - (-3) = 3 next terms are 9 + 3 = 12, 12 + 3 = 15 next 2 terms = 12, 15

OpenStudy (taylorterminate):

If u didnt understand: https://answers.yahoo.com/question/index?qid=20110117220735AAZHyhk

OpenStudy (taylorterminate):

@sheesh28